# Low-Rank Interconnected Adaptation across Layers

### v2

@@ 4.5 Understanding Lily
Does Lily Exhibit Selectivity? Lily uses routers to assign varying weights to different HP experts, thereby achieving selective information combination. We illustrate this selectivity in Fig. 6. We use a setup with three HP experts and select three layer levels (1, 13, 22) to calculate the total weight assigned to each expert. The results reveal a clear selectivity: for different layers, the router assigns significantly different weights to different HP experts. For instance, on Cifar100, the middle layer is predominantly dominated by HP 2, whereas the deep layer is primarily dominated by HP 1 and HP 2. In contrast, on Retinopathy, both the middle and deep layers are dominated by HP 3. This selectivity ensures that, even when different layers share information, the inherent differences between layers are still taken into account, making the adaptation more flexible and comprehensive.
@@ 5 Conclusion
In this paper, we propose low-rank interconnected adaptation (Lily), a novel framework for efficient fine-tuning via inter-connectivity of adapters. Lily enables each layer to access information from others during adaptation through a hierarchical structure. Additionally, it successfully overcome the low-rank update limitation of LoRA, enabling high-rank update and therefore better adaptation capability. Our approach consistently improves performance across various modalities, model sizes, and architectures, surpassing existing methods with enhanced efficiency. In summary, Lily’s versatility and efficiency make it a promising approach for a wide range of applications.
@@ A.1 Model Structure and Design Intuition of Lily
Within the overall framework of Lily, we delve into specific implementation details and model design insights. First, we have established the relationship between LPs and HP: LP is confined to specific levels of layers, capturing features that enable the router to selectively assign weights to the HP experts. In contrast, HP is a model-wide module comprising multiple experts, each of which contains information from a particular level of layers. We highlight several key aspects which are not heavily discussed in the methodology section:
A.1.1 Number of LPs
Since LP is limited to specific layers, the simplest approach would be to place an LP at each layer of the module to be adapted (e.g., the query transformation in MHSA). However, this setup may not be necessary: the importance of each layer varies, and many layers have significantly lower importance than others [Zhang et al., 2023]. To achieve greater parameter efficiency, we can set up fewer LPs, with each LP focusing on a level of layers rather than a single layer. For example, an LP can focus on shallow layers (e.g., layers 0, 1, 2, etc.) or deep layers. To enable a single LP to handle multiple layers, we can share an LP across multiple layers. By doing so, we eliminate the redundancy of having an LP at each layer, reduce the number of parameters, and increase efficiency. This is exactly the strategy adopted by most of the experiments.
A.1.2 Number of HP Experts
Regarding HP, the number of experts can be arbitrarily set, enabling more flexible configurations. In our experiments, for the sake of simplicity, we set the number of HP experts equal to the number of LPs, thereby equating the granularity of LP and HP.
@@ A.1.3 Routers Setup
There are also different settings that can be employed for the router. First, we can bind the router to HP, resulting in only one router per model. However, since the number of parameters in the router is relatively small, having only one router per model may not lead to significant selectivity. Therefore, we can also bind the router to LP, configuring a separate router for each LP. Most of our experiments use the latter setup, but in the vision experiments on Vim, we use the single-router and no-lp-sharing setup to verify its effectiveness. From the results, we can see that this setup also performs well. As future work, we can verify the effectiveness of using the latter setup on Vim, which may potentially lead to superior performance.
@@ A.1.4 Hyperparameters
We detail the hyperparameters used in Lily. Specifically, we use lily_r to represent the hidden-dimension of the projectors: LPs and HPs. It serves the same function as r in LoRA. We use lily_s to represent the scaling factor used by Lily. It is mostly searched within the range of {0.01, 0.1, 1.0, 10.0, 100.0}. We use ne_1 to represent the number of LPs used in the model. Since the LPs can be shared as discussed in the previous section, ne_1 does not need to equal the number of layers in the model. We use ne_2 to represent the number of HP experts in the model-wide HP module. In our experiments, we set ne_1 = ne_2 to improve parameter-efficiency and simplicity.
@@ A.1.5 Design Intuition
Lily employs a hierarchical structure to enable updates with higher-ranks than LoRA. However, simply equally connecting all the HPs to the LPs can not achieve the best performance. From the perspective of feature and information utilization across layers, simply aggregating all HPs for an LP ignores the distinctness of the features from current layers. Meanwhile, it reduced the variability of the combinations of gradient projection matrices ($S_{i}$ and $C_{i,j}$ are constants now), making the rank of the weight update higher than that of LoRA (since multiple distinct random matrices are used), but still not high enough for the best performance because of the lack of variability during combination. Therefore, we introduce selectivity into the inter-connectivity as well as discussed below, making the combination of HPs data-dependent so that each $S_{i}$ is unique across the time-steps, enabling updates with even higher ranks. The approach has similarity to Hao et al. [2024], where random matrix is constantly resampled to ensure updates with higher ranks. We further conduct an analysis in Appendix G.
@@ A.2 Efficient Implementation for Weighted Combination
One intuitive implementation of the weighted combination in Lily is Eq. 17, from which we could observe that it perform $N_{e}$ times of matrix multiplication, $N_{e}$ times of scalar multiplication and $N_{e}$ times of matrix addition. Therefore, despite its intuitive nature, the computational burden of this approach is quite formidable.
However, Eq. 18 which is adopted in Lily only utilize $N_{e}$ times of scalar multiplication, $N_{e}$ times of matrix addition and 1 time of matrix multiplication. This saves roughly $N_{e}$ times of matrix multiplication, which can be significant as the size of the model and number of adaptation targets increases. For a $x^{\prime}$ size of $\mathbb{R}^{N\times d}$ and a $P_{H}\in\mathbb{R}^{d\times C}$, the floating-point operation (FLOPs) of these two implementation are:
FLOPs $=\sum_{1}^{N_{e}}(2NdC)+\sum_{1}^{N_{e}}(dC)+\sum_{1}^{N_{e}}(NC)$ (21)
$=N_{e}\times(2NdC+dC+NC)$ (Intuitive)
@@
$=2dC\times(N+N_{e})$ (Lily)
from which we can easily observe that the approach adopted by Lily requires less computation and therefore provides more speed and efficiency during the fine-tuning process. Under the setting of $N=1024,d=16,C=768,N_{e}=4$, the FLOPs of the intuitive approach would be $0.104$ GFLOPs while in Lily it is merely $0.025$ GFLOPs, which could potentially lead to a 4X increase in speed.
@@ A.3 Actual implementation of Lily
Figure 7: Implementation of Lily in VTAB-1K benchmark.
We present the actual implementation of Lily in Fig. 7. For the example here, we choose the implementation from visual adaptation tasks (i.e., VTAB-1K benchmark). For LLM, the implementation is a bit more complicated because of modifications to the huggingface PEFT library [Mangrulkar et al., 2022], but the fundamental adaptation process is the same. Specifically, given an input, we first use the corresponding LP of the current layer to project it to a low-dimensional representation. After that, we use the low-dimensional representation to selectively assign weights for the HP experts. Once we obtain all the weights for the experts, we set out to combine these HP experts accordingly, as discussed in Appendix A.2. After the weighted combination, we use the obtained combined HP to project the low-dimensional representation to high-dimension, therefore acquiring the extra knowledge gained through adaptation.
@@ Appendix B Experimental settings
B.1 Hyper-parameters
A detailed description of the hyper-parameters used in Lily is in Appendix A.1.
@@ B.1.1 Commonsense Reasoning
The hyper-parameters used in commonsense reasoning experiments for MiLoRA, PiSSA are provided in Table 7, 6. Settings for Lily and LoRA using Falcon-Mamba as the backbone are provided in Table. 9 and 8. It can be noticed that Lily achieves the best performance by merely adapting the multi-head self attention module (MHSA) in LLaMA3-8B, while other compared methods adapt all the modules including MLP. Meanwhile, Lily employs the least amount of parameters, showcasing its excellent adaptation at low parameter-budget scenarios.
@@
Placement query, key, value, output, gate, MLP up, MLP down
@@ B.1.2 Natural Language Understanding
Specific hyper-parameter settings of Lily on GLUE benchmark are provided in Table 10. We fix the learning rate of both the backbone and the head as 5E-3 and tune the scaling factor lily_s $\in\{0.01,0.1,1.0\}$ instead. For the rank $r$ we fix it to 32 and the seed to 0.
@@
B.1.3 Visual Adaptation Benchmark
We provide the hyper-parameter for Lily on VTAB-1K benchmark in Table 11. Specifically, we fix the learning rate at 1E-3 with a weight decay of 1E-4. For ViT, we tune the scaling factor lily_s $\in\{0.01,0.1,1.0,10.0\}$ to maximize the performance, following Jie et al. [2023], Jie and Deng [2023]. For Vim, we fix lily_s to 1.0. Additionally, we search for the hyper-parameters ne_1 and ne_2 within the range {2, 3, 4}, as these numbers can divide the number of layers in the ViT model (12 in ViT-B). For vim, we use the implementation discussed in section A.1, which does not share LPs across layers. Therefore, ne_1 in this setting is fixed to number of layers in Vim (22 in this case), while we search ne_2 in {3,6} and {5,6,17} separately for Lily-S and Lily-L. Note that ne is only set for input projection in Vim. For delta transformation, we only use a single HP expert to reduce the parameter cost. In the experiments of ViT, the rank $r$ is fixed at 16. Meanwhile in Vim’s setting, we tune the ranks $r$ for the delta transformation module and the input projection module separately. We use $4,4$ and $4,8$ separately for Lily-S and Lily-L.
@@ B.2.3 Visual Adaptation Benchmark
Appendix C Does Sharing LP Results in Inferior Performance?
As mentioned earlier, we adopted a strategy of sharing the LP across most of our experiments, ensuring that the number of LP and HP experts is consistent. This approach offers two benefits: simplicity and enhanced parameter efficiency. By sharing the LP, we eliminate the need to set a separate LP for each layer, thereby reducing the overall parameter count.
Our decision to share the LP is based on the observation of overall redundancy among layers. Specifically, different layers have varying levels of importance [Zhang et al., 2023], and some less important layers do not require a dedicated LP. By not setting a separate LP for these layers, we avoid introducing extra parameter overhead while having a negligible impact on performance. To test that whether sharing LP results in inferior performance, we conduct experiments with no LP sharing on VTAB-1K. The results are shown in Table 15, from which we can observe that the best overall performance ($77.3\%$) is the same as that in the LP-sharing setting. This indicates that even if we employ one LP for each layer, the performance gain is negligible and many of the parameters are actually redundant. However, not sharing LPs results in extra parameter overhead and damages the parameter-efficiency of Lily. Therefore, LP-sharing is a great strategy to eliminate redundancy among LPs and boost the parameter-efficiency of Lily.
@@ Appendix D Where to Apply Lily in Transformers?
Table 15: Performance on VTAB-1K benchmark when applying Lily to various modules in Transformer. The implementation here does not share LP for simplicity (i.e., each layer has one LP).
@@
PEFT methods have been predominantly explored on the Transformer architecture, which consists of multi-head self-attention (MHSA) and multi-layer perceptron (MLP) as its core modules. In this section, we analyze the impact of fine-tuned modules on performance using Lily. Specifically, we compare Lily’s performance on the VTAB-1K benchmark under four settings:
@@
  • Applying Lily solely to the MLP module (denoted as "mlp").
  • Applying Lily to both the query and value transformation module in MHSA and the MLP module (denoted as "qvmlp").
  • Applying Lily to both the key and value transformation module in MHSA and the MLP module (denoted as "kvmlp").
To ensure a fair comparison, we tune the hyperparameters to maintain a similar parameter count across all settings. Additionally, to further investigate whether sharing the low-rank projection (LP) affects performance, we do not share the LP in this experiment. The results are presented in Table 15. We observe that the "kvmlp" setting achieves the best performance, with an average accuracy of $77.3\%$. In contrast, adapting only the MHSA module ("qv") yields the worst performance. Furthermore, we note that adapting both the MHSA and MLP modules (qvmlp and kvmlp) generally leads to superior results compared to adapting only one specific module (qv and mlp). This suggests that both MLP and MHSA play crucial roles in the overall model performance, and adapting both is essential for effective adaptation.
Notably, even when applying Lily solely to the MHSA module, which results in the worst performance among the four settings ($76.9\%$), it still outperforms LoRA by a significant margin ($0.5\%$). This underscores the efficiency of Lily, as it uses fewer parameters than LoRA even without LP sharing.
Appendix E Where to Apply Lily in Mamba?
Nearly all previous PEFT method studies have been centered around Transformers, while Mamba is a relatively new architecture, so there has been little research on PEFT methods on Mamba. In this section, we briefly analyze the pros and cons of adapting Mamba’s modules. In brief, a Mamba block consists of regular linear projection layers and a core component SSM module [Gu and Dao, 2023] [Zhu et al., 2024]. Specifically in SSM module, Mamba utilize parameters ($\Delta$, $A$, $B$, $C$) to transform an input sequence $x(t)$ to an output sequence $y(t)$ using a hidden state $h(t)$. The discretization process converts $A$ and $B$ into $\bar{A}$ and $\bar{B}$, respectively, using the time step size parameter $\Delta$. Structured state space models, inspired by continuous systems, can be computed similarly to RNNs or in the form of global convolution due to their linear time invariance (LTI) property. Mamba introduces a selective property to structured state space model, tying parameters to the current input, which breaks the LTI property and hinders parallel training. To address this, Mamba employs a hardware-aware algorithm, enabling its SSM module to possess the selective property and perform parallel training. To be specific, the discretization process can be expressed as:
$\bar{A}$ $=exp(\Delta A)$ (22)
$\bar{B}$ $=(\Delta A)^{-1}(exp(\Delta A)-I)\cdot\Delta B$

### v3

@@ 4.5.3 Does It Exhibit Selectivity?
Lily uses routers to assign varying weights to different $B$ experts, thereby achieving selective information combination. We illustrate this selectivity in Fig. 4. We use a setup with three $B$ experts and select three layer levels (1, 13, 22) to calculate the total weight assigned to each expert. The results reveal clear selectivity: for different layers, the router assigns significantly different weights to the $B$ experts. For instance, on Cifar100, the middle layer is predominantly dominated by $B$ 2, whereas the deep layer is primarily dominated by $B$ 1 and $B$ 2. In contrast, on Retinopathy, both the middle and deep layers are dominated by $B$ 3. This selectivity ensures that, even when different layers share information, the inherent differences between layers are still taken into account, making the adaptation more flexible and comprehensive.
4.5.4 What’s the Hardware Efficiency?
The dynamics of Lily obviously introduce complexity to the design of LoRA. In this section, we analyze how this affects the hardware efficiency of Lily compared to LoRA. We use the COLA task from the NLU experiments with RoBERTa-Base and run for 10 epochs. Additionally, we also report the runtime and GPU memory consumption in the Falcon-Mamba experiment.
The results are shown in Fig. 6, from which we can observe that the hardware efficiency of Lily is comparable to LoRA. Specifically, Lily slightly underperforms LoRA in the NLU experiment but performs on par with LoRA in the LLM experiment. In general, the introduced complexity of Lily does not prevent it from being an more effective PEFT method that is also hardware-friendly.
@@ 5 Conclusion
In this paper, we propose Low-Rank Interconnected Adaptation (Lily), a novel framework for efficient fine-tuning via the interconnectivity of adapters. Lily enables each layer to access information from others during adaptation through a hierarchical structure. Additionally, it successfully overcomes the low-rank update limitation of LoRA, enabling high-rank updates and, therefore, better adaptation capability under the same parameter budget. Our approach consistently improves performance across various modalities, model sizes, and architectures, surpassing existing methods while maintaining enhanced efficiency. In summary, Lily’s versatility and efficiency make it a promising approach for a wide range of applications.
6 Limitations
Although Lily has been experimentally evaluated in a wide range of scenarios, we have not explored all possible applications where PEFT could be used. These potential areas are left as directions for future work.
7 Ethics Statement
This work is an improvement upon LoRA. However, it could potentially be used for fine-tuning diffusion models or large language models (LLMs) for generating malicious content.
@@ A.1 Model Structure and Design Intuition of Lily
Within the overall framework of Lily, we delve into specific implementation details and model design insights. First, we establish the relationship between $A$ and $B$: $A$ is confined to specific levels of layers, capturing features that enable the router to selectively assign weights to the $B$ experts. In contrast, $B$ is a model-wide module comprising multiple experts, each of which contains information from a particular level of layers.
We highlight several key aspects that are not heavily discussed in the methodology section:
A.1.1 Number of $A$s
Since $A$ is limited to specific layers, the simplest approach would be to place an $A$ at each layer of the module to be adapted (e.g., the query transformation in MHSA). However, this setup may not be necessary, as the importance of each layer varies, and many layers have significantly lower importance than others (Zhang et al., 2023).
To achieve greater parameter efficiency, we can use fewer $A$s, with each $A$ focusing on a level of layers rather than a single layer. For example, an $A$ can focus on shallow layers (e.g., layers 0, 1, 2, etc.) or deep layers. To enable a single $A$ to handle multiple layers, we can share an $A$ across multiple layers. By doing so, we eliminate the redundancy of placing an $A$ at each layer, reduce the number of parameters, and improve efficiency.
This is exactly the strategy adopted in most of the experiments.
A.1.2 Number of $B$ Experts
Regarding $B$, the number of experts can be set arbitrarily, allowing for more flexible configurations. In our experiments, for the sake of simplicity, we set the number of $B$ experts equal to the number of $A$s, thereby equating the granularity of $A$ and $B$.
@@ A.1.3 Routers Setup
There are multiple possible configurations for the router. First, we can bind the router to $B$, resulting in only one router per model. However, since the number of parameters in the router is relatively small, having only one router per model may not provide significant selectivity. Therefore, we can also bind the router to $A$, configuring a separate router for each $A$.
Most of our experiments use the latter setup. However, in the vision experiments on Vim, we adopt the single-router and no-lp-sharing setup to evaluate its effectiveness. The results indicate that this setup also performs well.
As future work, we can further verify the effectiveness of the latter setup on Vim, which may potentially lead to superior performance.
@@ A.1.4 Hyperparameters
We detail the hyperparameters used in Lily. Specifically, we use Lily_r to represent the hidden dimension of the projectors: $A$s and $B$s. It serves the same function as $r$ in LoRA. We use Lily_s to represent the scaling factor used by Lily, which is primarily searched within the range {0.01, 0.1, 1.0, 10.0, 100.0}.
We use ne_1 to denote the number of $A$s used in the model. Since $A$s can be shared, as discussed in the previous section, ne_1 does not necessarily equal the number of layers in the model. Similarly, we use ne_2 to represent the number of $B$ experts in the model-wide $B$ module.
In our experiments, we set ne_1 = ne_2 to enhance parameter efficiency and maintain simplicity.
@@ A.1.5 Design Intuition
Lily employs a hierarchical structure to enable updates with higher ranks than LoRA. However, simply connecting all $B$s equally to $A$s does not yield the best performance. From the perspective of feature and information utilization across layers, merely aggregating all $B$s for an $A$ ignores the distinctiveness of features from the current layers. Meanwhile, this approach reduces the variability in the combinations of gradient projection matrices (since $S_{i}$ and $C_{i,j}$ become constants), making the rank of the weight update higher than that of LoRA (as multiple distinct random matrices are used), but still not high enough for optimal performance due to the lack of variability in the combination process.
To address this, we introduce selectivity into the interconnectivity, as discussed below, making the combination of $B$s data-dependent. This ensures that each $S_{i}$ is unique across time steps, enabling updates with even higher ranks. This approach is similar to that of Hao et al. (2024), where a random matrix is constantly resampled to maintain high-rank updates. We further analyze this in Appendix G.
@@ A.2 Efficient Implementation for Weighted Combination
A straightforward implementation of the weighted combination in Lily is to pass the inputs through all the experts and then sum the results. This approach requires $N_{e}$ matrix multiplications, $N_{e}$ scalar multiplications, and $N_{e}$ matrix additions. Despite its intuitive nature, the computational burden of this method is quite substantial.
However, Eq. 3, which is adopted in Lily, requires only $N_{e}$ scalar multiplications, $N_{e}$ matrix additions, and a single matrix multiplication. This optimization eliminates approximately $N_{e}$ matrix multiplications, which can significantly reduce computational costs as the model size and the number of adaptation targets increase.
For an input $x^{\prime}$ of size $\mathbb{R}^{N\times d}$ and a projection matrix $P_{H}\in\mathbb{R}^{d\times C}$, the floating-point operations (FLOPs) of these two implementations are:
FLOPs $=\sum_{i=1}^{N_{e}}(2NdC)+\sum_{i=1}^{N_{e}}(dC)+\sum_{i=1}^{N_{e}}(NC)$ (5)
$=N_{e}\times(2NdC+dC+NC),$
@@
$=2dC\times(N+N_{e}),$
From this, we can easily observe that the approach adopted by Lily requires fewer computations, thereby improving both speed and efficiency during the fine-tuning process. Under the setting of $N=1024,d=16,C=768,N_{e}=4$, the FLOPs for the intuitive approach amount to $0.104$ GFLOPs, whereas for Lily, it is merely $0.025$ GFLOPs, potentially leading to a 4×4$\times$4 × speedup.
@@ A.3 Actual Implementation of Lily
Figure 7: Implementation of Lily in the VTAB-1K benchmark.
We present the actual implementation of Lily in Fig. 7. In this example, we showcase its implementation for visual adaptation tasks, specifically in the VTAB-1K benchmark. For large language models (LLMs), the implementation is slightly more complex due to modifications to the Hugging Face PEFT library (Mangrulkar et al., 2022), but the fundamental adaptation process remains the same.
Specifically, given an input, we first use the corresponding $A$ of the current layer to project it into a low-dimensional representation. This low-dimensional representation is then used to selectively assign weights to the $B$ experts. Once the weights for all experts are obtained, we proceed to combine these $B$ experts accordingly, as discussed in Appendix A.2. After the weighted combination, the combined $B$ is used to project the low-dimensional representation back into a high-dimensional space, thereby incorporating the additional knowledge gained through adaptation.
@@ Appendix B Experimental Settings
B.1 Hyperparameters
A detailed description of the hyperparameters used in Lily is provided in Appendix A.1.
@@ B.1.1 Commonsense Reasoning
The hyperparameters used in commonsense reasoning experiments for MiLoRA and PiSSA are provided in Tables 7 and 6. The settings for Lily and LoRA using Falcon-Mamba as the backbone are presented in Tables 9 and 8.
[…]
@@
[…]
@@ B.1.2 Natural Language Understanding
The specific hyperparameter settings for Lily on the GLUE benchmark are provided in Table 10. We fix the learning rate of both the backbone and the head at $5\times 10^{-3}$ and instead tune the scaling factor Lily_s, where Lily_s $\in\{0.01,0.1,1.0\}$. The rank $r$ is fixed at 32, and the random seed is set to 0. The baseline results are taken from FourierFT (Gao et al., 2024b).
@@
B.1.3 Visual Task Adaptation Benchmark
We provide the hyperparameters for Lily on the VTAB-1K benchmark in Table 11. Specifically, we fix the learning rate at $1\times 10^{-3}$ with a weight decay of $1\times 10^{-4}$. For ViT, we tune the scaling factor Lily_s $\in\{0.01,0.1,1.0,10.0\}$ to maximize performance, following Jie et al. (2023) and Jie and Deng (2023). For Vim, we fix Lily_s to 1.0. Additionally, we search for the hyperparameters ne_1 and ne_2 within the range {2, 3, 4}, as these numbers divide the number of layers in the ViT model (12 in ViT-B).
For Vim, we use the implementation discussed in Section A.1, which does not share $A$s across layers. Therefore, ne_1 in this setting is fixed to the number of layers in Vim (22 in this case), while we search for ne_2 in {3, 6} and {5, 6, 17} separately for Lily-S and Lily-L. Note that ne is only set for the input projection in Vim. For the delta transformation, we use only a single $B$ expert to reduce the parameter cost.
In the ViT experiments, the rank $r$ is fixed at 16. Meanwhile, in Vim’s setting, we tune the ranks $r$ for the delta transformation module and the input projection module separately. We use $(4,4)$ and $(4,8)$ separately for Lily-S and Lily-L.
@@ B.2.3 Visual Adaptation Benchmark
Appendix C Does Sharing $A$s Result in Inferior Performance?
As mentioned earlier, we adopted a strategy of sharing the $A$ across most of our experiments, ensuring that the number of $A$s and $B$ experts is consistent. This approach offers two key benefits: simplicity and enhanced parameter efficiency. By sharing the $A$, we eliminate the need to set a separate $A$ for each layer, thereby reducing the overall parameter count.
Our decision to share the $A$ is based on the observation of overall redundancy among layers. Specifically, different layers have varying levels of importance (Zhang et al., 2023), and some less important layers do not require a dedicated $A$. By not setting a separate $A$ for these layers, we avoid introducing unnecessary parameter overhead while maintaining negligible impact on performance. To test whether sharing $A$ results in inferior performance, we conducted experiments without $A$ sharing on the VTAB-1K benchmark. The results, shown in Table 15, indicate that the best overall performance ($77.3\%$) is the same as in the $A$-sharing setting. This suggests that even when we employ one $A$ for each layer, the performance gain is negligible, and many of the parameters are, in fact, redundant. However, not sharing $A$s leads to additional parameter overhead, which reduces the parameter efficiency of Lily. Therefore, $A$-sharing is an effective strategy to eliminate redundancy among $A$s and enhance the parameter efficiency of Lily.
@@ Appendix D Where to Apply Lily in Transformers?
Table 15: Performance on VTAB-1K benchmark when applying Lily to various modules in Transformer. The implementation here does not share $A$ for simplicity (i.e., each layer has one $A$).
@@
Appendix E Performance Analysis on VTAB-1K Benchmark with Lily on Transformer Modules
[…]
@@ Appendix E Performance Analysis on VTAB-1K Benchmark with Lily on Transformer Modules
[…]
To ensure a fair comparison, we tune the hyperparameters to maintain a similar parameter count across all settings. Additionally, to further investigate whether sharing the low-rank projection ($A$) affects performance, we do not share $A$ in this experiment.
[…]
Notably, even when applying Lily solely to the MHSA module, which results in the worst performance among the four settings ($76.9\%$), it still outperforms LoRA by a significant margin ($0.5\%$). This underscores the efficiency of Lily, as it uses fewer parameters than LoRA, even without $A$ sharing.
Appendix F Where to Apply Lily in Mamba?
Nearly all previous PEFT method studies have focused on Transformers, while Mamba is a relatively new architecture, and therefore, there has been little research on PEFT methods for Mamba. In this section, we briefly analyze the pros and cons of adapting Mamba’s modules. A Mamba block consists of regular linear projection layers and a core component, the SSM module (Gu and Dao, 2023), (Zhu et al., 2024). Specifically, in the SSM module, Mamba utilizes parameters ($\Delta$, $A$, $B$, $C$) to transform an input sequence $x(t)$ into an output sequence $y(t)$ using a hidden state $h(t)$. The discretization process converts $A$ and $B$ into $\bar{A}$ and $\bar{B}$, respectively, using the time step size parameter $\Delta$. Structured state space models, inspired by continuous systems, can be computed similarly to RNNs or in the form of global convolution due to their linear time invariance (LTI) property. Mamba introduces a selective property to the structured state space model, tying parameters to the current input. This breaks the LTI property and hinders parallel training. To address this, Mamba employs a hardware-aware algorithm, enabling its SSM module to possess the selective property while performing parallel training.
To be specific, the discretization process can be expressed as:
$\bar{A}$ $=\exp(\Delta A)$ (6)
$\bar{B}$ $=(\Delta A)^{-1}(\exp(\Delta A)-I)\cdot\Delta B$